# Dynamic Relative Representations for Goal-Oriented Semantic Communications

### v2

@@ I Introduction
Classic communication paradigms prioritize the accurate transmission of individual symbols or bits without explicit consideration of their underlying meaning. As Shannon identified, communication can be understood through three levels: the syntactic level (focusing on the reliable transmission of signals), the semantic level (concerned with the meaning conveyed by those signals), and the effectiveness level (ensuring the message achieves its intended purpose) [1]. However, the exponential growth of data-intensive wireless applications demands a fundamental shift in this approach. The sheer volume of information required by applications like autonomous vehicles and augmented reality overwhelms the capacity of traditional bit-centric communication systems. In contrast, modern networks increasingly demand the exchange of information with shared meaning [2]. Semantic communication (SemCom) addresses this need by embedding meaning directly into transmissions, moving beyond the mere transmission of bits [3, 4, 5, 6]. To achieve this, SemCom relies heavily on artificial intelligence (AI) and deep neural networks (DNNs) to extract and transmit key semantic features from data rather than the raw data itself. Instead of focusing on the perfect reconstruction of raw symbols, SemCom prioritizes transmitting the underlying meaning, offering inherent robustness to noise and imperfect transmission conditions. At the receiver side, a DNN decoder reconstructs and interprets the intended message. The flexibility and resilience of SemCom make it ideal for emerging applications where devices operate in dynamic and resource-constrained environments [7]. Building on this idea, novel approaches to semantic joint source and channel coding [5], semantic extraction and compression [8, 9], goal-oriented system design and optimization [10, 7], semantic reasoning [11], and semantic communications based on generative AI [12] have been proposed in the literature. Despite these advancements, disruptions caused by misinterpretations of semantic information, known as semantic noise, pose a significant challenge. Such disturbances often arise from discrepancies in how different devices represent and interpret data, leading to errors in understanding, even when the physical transmission is error-free. This issue is particularly acute in dynamic edge environments, where various devices are involved, and complex system-wide retraining is not feasible. Misalignments in high-dimensional data representations can severely disrupt semantic understanding under these conditions [13]. This underscores the need for mechanisms that mitigate semantic mismatches without necessitating system updates.
@@
II Relative representation for Semantic Communication
SemCom aims to exchange the meaning conveyed by data rather than just the raw data itself. To achieve this, transmitters (TXs) and receivers (RXs) typically employ DNNs to encode and decode semantic information, transmitting absolute representations (raw DNN encoder outputs), as illustrated in Figure 1. In this scenario, semantic noise can arise from discrepancies in how different encoders semantically represent the same information due to stochastic factors or design choices. These misalignments introduce misinterpretations at the decoder, disrupting the overall communication objective.
To address this, we introduce RelRep into the SemCom framework, enabling zero-shot stitching – the ability to communicate effectively across diverse encoders without the need for retraining. With RelRep, a TX can switch encoders on the fly, and the RX can still interpret the message without needing to retrain itself for each new encoder. The rationale of RelRep for SemCom is explained in the sequel.
@@ III-A System Model
We assume a simplified RX that incurs negligible decoding delay. Perfect channel state information (CSI) is assumed to be available at the TX. We work with a time-slot-based model, where the system adaptively and dynamically optimizes resource allocation within each time-slot. In the following, we will describe our system model in detail .
@@ III-C Algorithmic Solution via Stochastic Optimization
$R_{t}^{*}$ $=\left[\frac{2B}{\log 2}W\Bigg{(}\frac{1}{2}\sqrt{\frac{(Z_{t}+Y_{t}){n_{{\mathbb{A}}_{t}}}q|h_{t}|^{2}\ln 2}{BVN_{0}}}\Bigg{)}\right]_{R^{MIN}}^{R^{MAX}}$ (11)

### v3

@@ I Introduction
Classic communication paradigms prioritize the accurate transmission of individual symbols or bits without explicit consideration of their underlying meaning. As Weaver identified, communication can be understood through three levels: the syntactic level (focusing on the reliable transmission of signals), the semantic level (concerned with the meaning conveyed by those signals), and the effectiveness level (ensuring the message achieves its intended purpose) [1]. However, the exponential growth of data-intensive wireless applications demands a fundamental shift in this approach. The sheer volume of information required by applications like autonomous vehicles and augmented reality overwhelms the capacity of traditional bit-centric communication systems. In contrast, modern networks increasingly demand the exchange of information with shared meaning [2]. Semantic communication (SemCom) addresses this need by embedding meaning directly into transmissions, moving beyond the mere transmission of bits [3, 4, 5, 6]. To achieve this, SemCom relies heavily on artificial intelligence (AI) and deep neural networks (DNNs) to extract and transmit key semantic features from data rather than the raw data itself. Instead of focusing on the perfect reconstruction of raw symbols, SemCom prioritizes transmitting the underlying meaning, offering inherent robustness to noise and imperfect transmission conditions. At the receiver side, a DNN decoder reconstructs and interprets the intended message. The flexibility and resilience of SemCom make it ideal for emerging applications where devices operate in dynamic and resource-constrained environments [7]. Building on this idea, novel approaches to semantic joint source and channel coding [5], semantic extraction and compression [8, 9], goal-oriented system design and optimization [10, 7], semantic reasoning [11], and semantic communications based on generative AI [12] have been proposed in the literature. Despite these advancements, disruptions caused by misinterpretations of semantic information, known as semantic noise, pose a significant challenge. Such disturbances often arise from discrepancies in how different devices represent and interpret data, leading to errors in understanding, even when the physical transmission is error-free. This issue is particularly acute in dynamic edge environments, where various devices are involved, and complex system-wide retraining is not feasible. Misalignments in high-dimensional data representations can severely disrupt semantic understanding under these conditions [13]. This underscores the need for mechanisms that mitigate semantic mismatches without necessitating system updates.
@@
II Relative Representation Motivations
SemCom aims to exchange the meaning conveyed by data rather than just the raw data itself. To achieve this, transmitters (TXs) and receivers (RXs) typically employ DNNs to encode and decode semantic information, transmitting absolute representations (raw DNN encoder outputs). In this scenario, semantic noise can arise from discrepancies in how different encoders semantically represent the same information due to stochastic factors or design choices. These misalignments introduce misinterpretations at the decoder, disrupting the overall communication objective.
To address this, we introduce RelRep into the SemCom framework, enabling zero-shot stitching – the ability to communicate effectively across diverse encoders without the need for retraining. With RelRep, a TX can switch encoders on the fly, and the RX can still interpret the message without needing to retrain itself for each new encoder, as in Figure 1. The rationale of RelRep for SemCom is explained in the sequel.
@@ III-A System Model
We assume a simplified RX that incurs negligible decoding delay. Perfect channel state information (CSI) is assumed to be available at the TX. We work with a time-slot-based model, where the system adaptively and dynamically optimizes resource allocation within each time slot. In the following, we will describe our system model in detail .
@@ III-C Algorithmic Solution via Stochastic Optimization
$R_{t}^{*}$ $=\left[\frac{2B}{\ln 2}W\Bigg{(}\frac{1}{2}\sqrt{\frac{(Z_{t}+Y_{t}){n_{{\mathbb{A}}_{t}}}q|h_{t}|^{2}\ln 2}{BVN_{0}}}\Bigg{)}\right]_{R^{MIN}}^{R^{MAX}}$ (11)